# Multiwavelength coverage of the 2024 periastron passage of PSR B125963 / LS 2883

### v1

@@ Abstract
PSR B125963 is a gamma-ray binary system with a 48 ms radio pulsar orbiting around an O9.5Ve star, LS 2883, in a highly eccentric yr long orbit. Close to the periastron the system is detected from radio up to the TeV energies due to the interaction of LS 2883 and pulsar’s outflows. The observations of last 4 periastra passages taken in 2010-2021 demonstrate periastron to periastron variability at all wavelength, probably linked to the state of the decretion disk. In this paper we present the results of our optical, radio and X-ray observational campaigns on PSR B125963 performed in 2024 accompanied with the analysis of the publicly available GeV Fermi/LAT data. We show that this periastron passage was characterised by the early flaring of X-rays before the periastron passage and GeV emission after the periastron passage, which can be explained by a larger size of the decretion disk as supported by the optical observations. The structure of the GeV flare is also in agreement with the disruption of the large dense disk. The possible X-ray/radio correlation was observed only during the post-periastron rise of X-ray and radio emission.
@@
First discovered during the Parkes high-frequency radio survey in 1990 (Johnston et al., 1992), PSR B125963 is a classical gamma-ray loud binary. In this system a young 48 ms radio pulsar is orbiting around an O9.5Ve massive star, LS 2883, in a highly eccentric () 3.4 years long orbit (Negueruela et al., 2011; Shannon, Johnston & Manchester, 2014; Miller-Jones et al., 2018). The decretion disk of LS 2883 is inclined to the orbital plane by about 35∘ (Shannon, Johnston & Manchester, 2014) and the pulsed radio emission is eclipsed by the disk around the periastron. Unpulsed radio emission appears at about the time the pulsar approaches the disk for the first time, showing two maxima (peaks) during the first and second disk crossing, which greatly exceed the flux level of the pulsed emission. The position of these radio peaks roughly coincides with the peaks in the X-ray light curve (see e.g. Chernyakova et al., 2021).
@@
The observed radio to TeV emission from this system is generally believed to be the result of the interaction of the relativistic electrons of the pulsar wind with the LS 2883 radiation and wind outflow. The radio and X-ray emission is usually assumed to be a result of synchrotron emission (Tavani & Arons, 1997; Connors et al., 2002; Chernyakova et al., 2020), though that the emission region(s) and electron population(s) are still under debate. The TeV emission is due to the Inverse Compton (IC) scattering of the soft photons on the relativistic electrons of the pulsar wind (Kirk, Ball & Skjæraasen, 1999; Chernyakova et al., 2020). The origin of the GeV emission from this system is still under debate. The observed variability of the GeV emission on timescales as short as 15 minutes (Johnson et al., 2018) and the lack of correlation with any other energy band makes it difficult to attribute the observed GeV emission due to pure synchrotron or IC emission. Optical observations show an increase in the equivalent width of the H emission line around periastron, and a subsequent decrease starting around 30 days after periastron, which is linked to changes in the structure of the circumstellar disk (Chernyakova et al., 2017, 2021, 2024; van Soelen et al., 2016). A possible solution to the origin of the GeV flare was proposed by Chernyakova et al. (2020). In this paper it was shown that a combination of IC emission with the bremsstrahlung emission of the relativistic electrons drifting along the shock and interacting with the clumps produced due to the disruption of the disk can explain both the observed luminosity and the structure of the flare.
@@
Observations of PSR B125963 before the 2021 periastron allowed us to propose a relation between the GeV flare and the disruption of the disk, suggested by the decrease of the equivalent width of the H emission line (Chernyakova et al., 2020). Observations of PSR B125963 during its 2021 periastron passage, however, have clearly demonstrated that the evolution of the disk is not directly linked to the start of the GeV flare, as the evolution of the H equivalent width in 2014 and 2021 was very similar, while the start of GeV flare in 2021 was very delayed and began only 60 days after the periastron (Chernyakova et al., 2021). Observations of the 2021 periastron passage also demonstrated other new features, like the correlation of the radio/X-ray emission during the second X-ray peak (and the subsequent disappearance of this correlation afterwards), the sudden appearance of the third X-ray peak (Chernyakova et al., 2021, 2024), and evidence for X-ray/TeV correlation (H. E. S. S. Collaboration et al., 2024).
@@ 2.1 Radio Data
For the 2024 periastron, PSR B125963 was observed in the radio band with ATCA, under project code C3582. This observing campaign began approximately 20 days before periastron (2024 June 11), finishing 79 days after the periastron (2024 Sep 17). Data were recorded at two central frequencies, centred at 5.5 GHz and 9 GHz, with 2 GHz of bandwidth at each frequency (Wilson et al., 2011).
@@
The Common Astronomy Software Applications (CASA), version 6.6.3.22 (CASA Team et al., 2022) was used to analyse the data. When starting the analysis, the data had to be flagged by removing radio frequency interference (RFI) and systematic issues. The data were then calibrated following standard procedures in CASA.111https://casaguides.nrao.edu/index.php/Main_Page Imaging was carried out using tclean within CASA. The calibrated integrated flux value was found by fitting a 2-D Gaussian to the point source in the image with the imfit function.
@@ 2.2 Optical Data
Optical coverage of the 2024 periastron passage of PSR B125963 was carried out with the Robert Stobie Spectrograph (RSS; Burgh et al., 2003) on the Southern African Large Telescope (SALT; Buckley, Swart & Meiring, 2006). Observations were obtained from 60 d before (2024 May 1) until d after periastron (2024 Aug 5) in order to trace the behaviour of the H emission line (produced by the circumstellar disk) as the pulsar passed through periastron. All observations were taken with the PG2300 grating, with the 0.6″ long slit (achieving a resolving power of 11 000), and consisted of 1530 s exposures. The data reduction and spectral extraction followed the same procedure as discussed in Chernyakova et al. (2021, 2024). All the data was first pre-reduced by the SALT pipeline, and then flat correction, wavelength calibration and spectral extraction was performed, and each night’s observations were averaged together, continuum corrected with a low-order polynomical, and barycentric corrected, using a python script relying on numpy, scipy, astropy and pyraf. The equivalent width (EW) of the H emission line was then measured by numerically integrating over the wavelength range Å, and the uncertainty was estimated following Vollmann & Eversberg (2006), with the signal-to-noise ratio estimated as in Stoehr et al. (2008). The variation in H EW for the 2014 - 2024 periastra is shown in Fig. 1.
@@ 2.3.2 NICER
We used data from the monitoring campaign of PSR B125963, where spectral extraction was possible and the spectrum was not dominated by the background noise, consisting of 27 observations covering MJD 60472–60539 (ObsIDs 5203610101-6, 8-15, 17-21, 27, 31-32, 35-36). NICER data were reduced using the nicerdas software and CALDB version 20240206 according to the official data analysis threads.333https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/ The data were processed using the nicerl2 task, then the cleaned event files were used to generate spectral products using the nicerl3-spect task, incorporating the SCORPEON background model. For each observation an “orbit-night” spectrum was extracted using standard threshold setting. If a “night” spectrum had zero effective exposure, a “day” spectrum was extracted instead. NICER spectra were fitted using a simple absorbed power-law. In accordance with the official SCORPEON analysis manual, noise_norm, lhb_em and gal_nh parameters of SCORPEON model were frozen.444https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/scorpeon-xspec/ The PSR B125963 spectra were approximated using simple absorbed power-law using “cstat” statistic.
@@ 2.4 GeV Data
The adapted model of the region included templates for the Galactic and isotropic diffuse emission components provided by the Fermi/LAT collaboration, as well as all sources from the latest Fermi/LAT 4FGL-DR4 catalogue (Abdollahi et al., 2020), with the spectral templates selected according to the catalogue. At the initial stage of the analysis we assumed all spectral parameters of sources within around PSR B125963 to be free parameters and fixed all spectral parameters of sources between and to their catalogue values. We performed the fitting of the described model to the whole available data-set. For the subsequent analysis, we fixed all (except normalisations) free spectral parameters to their best-fit values, and removed all of the weak sources detected with test-statistics from the model.
@@
The obtained light curves suggest the enhancement of the GeV emission during +19 to +77 days (MJD 60510 – 60568) after the periastron. Below we refer to this period as the “Fermi/LAT flare” during the 2024 periastron passage. To study the potential spectral-shape variability of the source during different periods of the periastron passage we built the spectrum of PSR B125963 for the periods , days around the periastron, as well as during the Fermi/LAT flare period. These periods were selected to be similar to those used by Chernyakova et al. (2021) which allows us to make a direct comparison between the current and 2021 periastron passages. The spectra of 2024 and 2021 periastra passages are shown in Figure 4.
We found that PSR B125963 is detected during all considered time periods with a test-statistics (TS) of for days; for days and for the period of the flare ( days). Due to a low detection significance of the source during days and days periods we fit the spectra during these periods with a powerlaw (PL) model, additionally fixing the powerlaw index for the days period to . For the flaring period the spectrum of PSR B125963 is significantly curved and can be described with a superexponential cutoff-powerlaw (ECSPL) model. The best-fit spectral parameters in the GeV band are summarized in Table 1.
@@ 3 Results and Discussion
The 2024 periastron passage of PSR B125963 demonstrated a peculiar behaviour of the system from radio to GeV energies. The H equivalent width (Figs. 1 & 2) showed a general behaviour which was similar to previous periastra: the EW was fairly constant before the time of the first disk crossing; it increased and showed a local maxima around the first disk crossing before decreasing; the EW then increased and peaked after the periastron, before decreasing again. However, there are notable differences this time. First, the increase around the first disk crossing was noticeably larger than it was in 2021, and second the first maxima in 2024 appears to be day earlier than in 2017. Unfortunately, the gaps in the data do not allow us to place a tighter constraint on this. Second, while the EW is similar up to periastron, it is significantly larger from approximately 20 days after periastron. This difference in the EW would be inline with the pulsar interacting earlier with the circumstellar disc, either because of a larger disc, or due to a small shift in the disk orientation.
@@
Such an assumption is further supported by the X-ray observations, see Fig. 2. The bright and narrow first X-ray maximum was observed at days, i.e., a few days earlier than ever observed before. The second X-ray peak is characterized by a sharp rise at days, in line with previous observations, following by a period of highly variable enhanced emission until at least days. Such a behavior can be understood if, during the second passage, the pulsar perturbed the disk and/or dragged a large amount of the decretion disk’s material along with it. In this case the apex of the pulsar-stellar wind interaction cone could have remained relatively close to the pulsar for a longer time, i.e. in a higher magnetic field region resulting in enhanced X-ray emission. The strong quasi-random X-ray variability at to days can be connected to large-scale density variations in the perturbed disk which push the tip of the emission cone closer to/further from the pulsar.
@@
Unfortunately our ATCA data does not cover the rise in X-ray emission during the first X-ray peak which, therefore, does not not allow us to study the radio/X-ray correlation during this period. We started our observations just before the first X-ray maximum, and the radio demonstrated a smooth decay until 11 days before the periastron showing no correlation with the X-ray data. The rapid radio increase after the periastron (at days) occurred with the rise of the X-ray emission, and closely matched the data from previous periastron passages. This time, however the X-ray/radio correlation disappeared right after the beginning of the X-ray flux decay days after the periastron, contrary to the behaviour observed during the 2021 periastron passage, where the correlation was seen until days after the periastron.
@@
The absence of a correlation can be connected to different population of electrons producing radio and X-ray emission as suggested by Chernyakova et al. (2021). The X-ray light curve allows two interpretation of the late-phases of the second X-ray peak: (i): a broad 2nd peak with occasional initial radio/X-ray coupling; (ii): the presence of the 3rd X-ray peak, similar to one observed in 2021 (Chernyakova et al., 2020), but appearing much earlier at days.
@@
The radio flux density in 2024 reached a historical maximum 16 days after the periastron (see e.g. compilations of historical radio data in Chernyakova et al., 2021, and references therein). This maximum was followed by a rapid decay during which the radio emission dropped to a level comparable with the level of the pulsed emission at days after the periastron, much earlier than previously observed. The indexes of the radio emission is presented in the fourth panel of Figure 2 (green circles) along with the X-ray indexes (blue points) and the slope between the radio to X-ray (magenta squares) . In 2024 the indexes of the radio data were much more chaotic than in 2021, where they were systematically softer than X-ray ones.
@@
The GeV-band emission demonstrated a sharp rise from days, which peaked at days, at a time comparable to the 2010 periastron and substantially earlier than observed for the 2021 periastron. After a gradual decay of the emission a second rise was observed at days. A number of short (up to 6 min duration) flares with energetics exceeding spin-down by up to factor of were observed during both periods of high GeV flux, see Fig. 3.
@@
The interaction of the pulsar wind with the outflow of LS2883 led to the formation of a shock cone directed away from the massive star, with the axis lying along the line connecting the pulsar and LS 2883 (see Figure 5 in Chernyakova et al., 2017)). In the model of Chernyakova et al. (2017) the GeV emission is explained as an emission of the relativistic pulsar wind electrons flowing along the shock with a moderate Lorentz factor . Similar to the hollow cone model described in Neronov & Chernyakova (2008) the GeV Fermi/LAT flare occurs when the pulsar wind/decretion disk interaction cone is oriented towards the observer. The short min-long sub-flares with energetics significantly exceeding spin-down luminosity are explained as bremsstrahlung radiation that occurs due to the entrance of the clumps of the decretion disk into the interaction cone.
@@
In 2024 the Fermi/LAT flare occurred close to the end of the second X-ray peak, during the period of quasi-random X-ray variability. As discussed above during this period the regular structure of the decretion disk was strongly perturbed by the pulsar. Still, the density of the material was sufficiently high to form a relatively narrow emission cone and the rise of the GeV emission at to days.
@@
The second rise of the GeV emission observed at to days can be connected to the eventual disruption of the disk and the production of a large number of clumps which reach the interaction cone. This hypothesis is further supported by a larger number of short sub-flares seen during the second rise of the GeV emission, comparable to the first one (see Figure 3). The Bayesian-blocks time binning light curve indicates the presence of 5 sub-flares with the mean flux exceeding the spin-down luminosity for the period to days and 7 sub-flares for the period to days. Such a behavior can be indicative of a larger number of clumps reaching the interaction cone when the pulsar was significantly far from the periastron.
@@
Interestingly, the Bayesian analysis revealed the presence of short timescale variability well before the periastron at and days, with fluxes comparable to the spin-down luminosity. While the Fermi/LAT sensitivity does not allow us to firmly conclude on the exact flux level of these flaring events we note that pre-periastron flares with spin-down-exceeding luminosity are hard to interpret within the discussed model. We also note that pre-periastron flares with a similar luminosity were not observed during 2010-2021 periastron passages.
@@
The spectral behavior of the system in the GeV band differs from the one observed during the 2017-2021 periastron passages. Contrary to previous periastra, in 2024, PSR B125963 is detected with a higher significance before the periastron ( to days, ) and a lower significance after the periastron ( to days, ). The low significance of the detection during these time periods did not allow us to detect a spectral curvature contrary to previous passages. The flaring-period-averaged flux ( to days) is also somewhat lower than the on seen in 2017 (see Fig. 4 and Table 1).

### v2

@@ Abstract
PSR B125963 is a gamma-ray binary system with a 48 ms radio pulsar orbiting around an O9.5Ve star, LS 2883, in a highly eccentric yr long orbit. Close to the periastron the system is detected from radio up to the TeV energies due to the interaction of the stellar wind from LS 2883 and the pulsar’s relativistic outflow. Observations of the last four periastron passages, taken in 2010-2021, demonstrate periastron-to-periastron variability at all wavelengths, probably linked to the state of the Be star’s decretion disk. In this paper we present the results of our optical, radio and X-ray observational campaigns on PSR B125963 performed in 2024 accompanied with the analysis of the publicly available GeV Fermi/LAT data. We show that this periastron passage was characterised by the early flaring of X-rays before the periastron passage and GeV emission after the periastron passage, which can be explained by a larger size of the decretion disk as supported by the optical observations. The structure of the GeV flare is also in agreement with the disruption of the large dense disk. The possible X-ray/radio correlation was observed only during the post-periastron rise of X-ray and radio emission.
@@
First discovered during the Parkes high-frequency radio survey in 1990 (Johnston et al., 1992), PSR B125963 is a classical gamma-ray loud binary. In this system a young 48 ms radio pulsar is orbiting around an O9.5Ve massive star, LS 2883, in a highly eccentric () 3.4 years long orbit (Negueruela et al., 2011; Shannon, Johnston & Manchester, 2014; Miller-Jones et al., 2018). The decretion disk of LS 2883 is inclined to the orbital plane by about 35∘ (Shannon, Johnston & Manchester, 2014) and the pulsed radio emission is eclipsed by the disk around periastron. Unpulsed radio emission appears at about the time the pulsar approaches the disk for the first time, showing two maxima (peaks) during the first and second disk crossing, which greatly exceed the flux level of the pulsed emission. The position of these radio peaks roughly coincides with the peaks in the X-ray light curve (see e.g. Chernyakova et al., 2021).
@@
The observed radio to TeV emission from this system is generally believed to be the result of the interaction of the relativistic electrons of the pulsar wind with the LS 2883 radiation and wind outflow. The radio and X-ray emission is usually assumed to be a result of synchrotron emission (Tavani & Arons, 1997; Connors et al., 2002; Chernyakova et al., 2020), though note that the emission region(s) and electron population(s) are still under debate. The TeV emission is due to the Inverse Compton (IC) scattering of the soft photons on the relativistic electrons of the pulsar wind (Kirk, Ball & Skjæraasen, 1999; Chernyakova et al., 2020). The origin of the GeV emission from this system is still under debate. The observed variability of the GeV emission on timescales as short as 15 minutes (Johnson et al., 2018) and the lack of correlation with any other energy band excludes one-zone models (e.g. Tam et al., 2011; Kong, Cheng & Huang, 2012) and makes it difficult to attribute the observed GeV emission to either synchrotron or IC emission (e.g. Khangulyan et al., 2012; Dubus & Cerutti, 2013; Tam et al., 2015; Chen et al., 2019). Optical observations show an increase in the equivalent width of the H emission line around periastron, and a subsequent decrease starting around 30 days after periastron, which is linked to changes in the structure of the circumstellar disk (Chernyakova et al., 2017, 2021, 2024; van Soelen et al., 2016). A possible solution to the origin of the GeV flare was proposed by Chernyakova et al. (2020). In this paper it was shown that a combination of IC emission with the bremsstrahlung emission of the relativistic electrons drifting along the shock and interacting with the clumps produced due to the disruption of the disk can explain both the observed luminosity and the structure of the flare.
@@
Observations of PSR B125963 before the 2021 periastron allowed us to propose a relation between the GeV flare and the disruption of the disk, suggested by the decrease of the equivalent width of the H emission line (Chernyakova et al., 2020). Observations of PSR B125963 during its 2021 periastron passage, however, have clearly demonstrated that the evolution of the disk is not directly linked to the start of the GeV flare, as the evolution of the H equivalent width in 2014 and 2021 was very similar, while the start of GeV flare in 2021 was very delayed and began only 60 days after periastron (Chernyakova et al., 2021). Observations of the 2021 periastron passage also demonstrated other new features, like the correlation of the radio/X-ray emission during the second X-ray peak (and the subsequent disappearance of this correlation afterwards), the sudden appearance of the third X-ray peak (Chernyakova et al., 2021, 2024), and evidence for X-ray/TeV correlation (H. E. S. S. Collaboration et al., 2024).
@@ 2.1 Radio Data
For the 2024 periastron, PSR B125963 was observed in the radio band with ATCA, under project code C3582. This observing campaign began approximately 20 days before periastron (2024 June 11), finishing 79 days after the periastron (2024 Sep 17). Data was recorded at two central frequencies, centred at 5.5 GHz and 9 GHz, with 2 GHz of bandwidth at each frequency (Wilson et al., 2011).
@@
The Common Astronomy Software Applications (CASA), version 6.6.3.22 (CASA Team et al., 2022) was used to analyse the data. When starting the analysis, the data had to be flagged by removing radio frequency interference (RFI) and systematic issues. The data were then calibrated following standard procedures in CASA.111https://casaguides.nrao.edu/index.php/Main_Page Imaging was carried out using tclean within CASA. The calibrated integrated flux value was found by fitting a 2-D Gaussian to the point source in the image with the imfit function, fixing the Gaussian full-width-at-half-maximum values to those of the synthesized beam.
@@ 2.2 Optical Data
Optical coverage of the 2024 periastron passage of PSR B125963 was carried out with the Robert Stobie Spectrograph (RSS; Burgh et al., 2003) on the Southern African Large Telescope (SALT; Buckley, Swart & Meiring, 2006). Observations were obtained from 60 d before (2024 May 1) until d after periastron (2024 Aug 5) in order to trace the behaviour of the H emission line (produced by the circumstellar disk) as the pulsar passed through periastron. All observations were taken with the PG2300 grating, with the 0.6″ long slit (achieving a resolving power of 11 000), and consisted of 1530 s exposures. The data reduction and spectral extraction followed the same procedure as discussed in Chernyakova et al. (2021, 2024). All the data was first pre-reduced by the SALT pipeline, and then flat correction, wavelength calibration and spectral extraction was performed, and each night’s observations were averaged together, continuum corrected with a low-order polynomial, and barycentric corrected, using a python script relying on numpy, scipy, astropy and pyraf. The equivalent width (EW) of the H emission line was then measured by numerically integrating over the wavelength range Å, and the uncertainty was estimated following Vollmann & Eversberg (2006), with the signal-to-noise ratio estimated as in Stoehr et al. (2008). The variation in H EW for the 2014 - 2024 periastra is shown in Fig. 1.
@@ 2.3.2 NICER
We used data from the monitoring campaign of PSR B125963, where spectral extraction was possible and the spectrum was not dominated by the background noise, consisting of 27 observations covering MJD 60472–60539 (ObsIDs 5203610101-6, 8-15, 17-21, 27, 31-32, 35-36). NICER data were reduced using the nicerdas software and CALDB version 20240206 according to the official data analysis threads.333https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/ The data were processed using the nicerl2 task, then the cleaned event files were used to generate spectral products using the nicerl3-spect task, incorporating the SCORPEON background model. For each observation an “orbit-night” spectrum was extracted using standard threshold settings. If a “night” spectrum had zero effective exposure, a “day” spectrum was extracted instead. NICER spectra were fitted using a simple absorbed power-law. In accordance with the official SCORPEON analysis manual, noise_norm, lhb_em and gal_nh parameters of SCORPEON model were frozen.444https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/scorpeon-xspec/ The PSR B125963 spectra were approximated using simple absorbed power-law using “cstat” statistic.
@@ 2.4 GeV Data
The adapted model of the region included templates for the Galactic and isotropic diffuse emission components provided by the Fermi/LAT collaboration, as well as all sources from the latest Fermi/LAT 4FGL-DR4 catalogue (Abdollahi et al., 2020), with the spectral templates selected according to the catalogue. At the initial stage of the analysis we assumed all spectral parameters of sources within around PSR B125963 to be free parameters and fixed all spectral parameters of sources between and to their catalogue values. We performed the fitting of the described model to all the available data. For the subsequent analysis, we fixed all (except normalisations) free spectral parameters to their best-fit values, and removed all of the weak sources detected with test-statistics from the model.
@@
The obtained light curves suggest the enhancement of the GeV emission during +19 to +77 days (MJD 60510 – 60568) after periastron. Below we refer to this period as the “Fermi/LAT flare” during the 2024 periastron passage. To study the potential spectral-shape variability of the source during different periods of the periastron passage we built the spectrum of PSR B125963 for the periods , days around periastron, as well as during the Fermi/LAT flare period. These periods were selected to be similar to those used by Chernyakova et al. (2021) which allows us to make a direct comparison between the current and 2021 periastron passages. The spectra of the 2024 and 2021 periastron passages are shown in Figure 4.
We found that PSR B125963 is detected during all considered time periods with a test-statistic (TS) of for days; for days and for the period of the flare ( days). Due to a low detection significance of the source during days and days periods we fit the spectra during these periods with a powerlaw (PL) model, additionally fixing the powerlaw index for the days period to . For the flaring period the spectrum of PSR B125963 is significantly curved and can be described with a superexponential cutoff-powerlaw (ECSPL) model. The best-fit spectral parameters in the GeV band are summarized in Table 1.
@@ 3 Results and Discussion
The 2024 periastron passage of PSR B125963 demonstrated a peculiar behaviour of the system from radio to GeV energies. The H equivalent width (Figs. 1 & 2) showed a general behaviour which was similar to previous periastra: the EW was fairly constant before the time of the first disk crossing; it increased and showed a local maximum around the first disk crossing before decreasing; the EW then increased and peaked after periastron, before decreasing again. However, there are notable differences this time. First, the increase around the first disk crossing was noticeably larger than it was in 2021, and second the first maximum in 2024 appears to be day earlier than in 2017. Unfortunately, the gaps in the data do not allow us to place a tighter constraint on this. Finally, the peak of EW after periastron is significantly larger than was ever observed before.
This difference in the EW would be inline with the pulsar interacting earlier with the circumstellar disc, either because of a larger disc, or due to a small shift in the disk orientation. The presence of the large disk up to is also indicated by infrared observations (Fujita et al., 2024).
@@
Such an assumption is further supported by the X-ray observations, see Fig. 2. The bright and narrow first X-ray maximum was observed at days, i.e., a few days earlier than ever observed before. The second X-ray peak is characterized by a sharp rise at days, in line with previous observations, followed by a period of highly variable enhanced emission until at least days. Such a behavior can be understood if, during the second passage, the pulsar perturbed the disk and/or dragged a large amount of the decretion disk’s material along with it. In this case the apex of the pulsar-stellar wind interaction cone could have remained relatively close to the pulsar for a longer time, i.e. in a higher magnetic field region resulting in enhanced X-ray emission. The strong quasi-random X-ray variability at to days can be connected to large-scale density variations in the perturbed disk which push the tip of the emission cone closer to/further from the pulsar.
@@
Unfortunately our ATCA data does not cover the rise in X-ray emission during the first X-ray peak which, therefore, does not not allow us to study the radio/X-ray correlation during this period. We started our observations just before the first X-ray maximum, and the radio demonstrated a smooth decay until 11 days before periastron. The rapid radio increase after periastron (at days) occurred with the rise of the X-ray emission, and closely matched the data from previous periastron passages. This time, however the X-ray/radio correlation disappeared right after the beginning of the X-ray flux decay days after the periastron, contrary to the behaviour observed during the 2021 periastron passage, where the correlation was seen until days after periastron.
@@
The absence of a correlation can be connected to different populations of electrons producing radio and X-ray emission as suggested by Chernyakova et al. (2021). X-ray light curve allows two interpretations of the late-phases of the second X-ray peak: (i): a broad 2nd peak with occasional initial radio/X-ray coupling; (ii): the presence of the 3rd X-ray peak, similar to one observed in 2021 (Chernyakova et al., 2020), but appearing much earlier at days.
@@
The radio flux density in 2024 reached a historical maximum 16 days after periastron (see e.g. compilations of historical radio data in Chernyakova et al., 2021, and references therein). This maximum was followed by a rapid decay during which the radio emission dropped to a level comparable with the level of the pulsed emission at days after the periastron, much earlier than previously observed. The indexes of the radio emission are presented in the fourth panel of Figure 2 (green circles) along with the X-ray indexes (blue points) and the slope between the radio to X-ray (magenta squares). In 2024 the indexes of the radio data were much more chaotic than in 2021, where they were systematically softer than the X-ray ones.
@@
The GeV-band emission demonstrated a sharp rise from days, which peaked at days, at a time comparable to the 2010 periastron and substantially earlier than observed for the 2021 periastron. Such an early rise of the GeV emission contradicts the precession disk model proposed by Chang et al. (2021) After a gradual decay of the emission a second rise was observed at to days. A number of short down to 6 min duration) flares with energetics exceeding the spin-down luminosity up to a factor of were observed during both periods of high GeV flux, see Fig. 3.
@@
The interaction of the pulsar wind with the outflow of LS2883 led to the formation of a shock cone directed away from the massive star, with the axis lying along the line connecting the pulsar and LS 2883 (see Figure 5 in Chernyakova et al., 2017)). In the model of Chernyakova et al. (2017) the GeV emission is explained as emission of the relativistic pulsar wind electrons flowing along the shock with a moderate Lorentz factor . Similar to the hollow cone model described in Neronov & Chernyakova (2008) the GeV Fermi/LAT flare occurs when the pulsar wind/decretion disk interaction cone is oriented towards the observer. The short min-long sub-flares with energetics significantly exceeding the spin-down luminosity are explained as bremsstrahlung radiation that occurs due to the entrance of the clumps of the decretion disk into the interaction cone.
@@
In 2024 the Fermi/LAT flare occurred close to the end of the second X-ray peak, during the period of quasi-random X-ray variability. As discussed above, during this period the regular structure of the decretion disk was strongly perturbed by the pulsar. Still, the density of the material was sufficiently high to form a relatively thin emission cone and the rise of the GeV emission at to days.
@@
The second rise of the GeV emission ( to days) can be connected to the eventual disruption of the disk and the production of a large number of clumps which reach the interaction cone. This hypothesis is further supported by a larger number of short sub-flares seen during the second rise of the GeV emission, comparable to the first one (see Figure 3). The Bayesian-blocks time binning light curve indicates the presence of 5 sub-flares with the mean flux exceeding the spin-down luminosity for the period to days and 7 sub-flares for the period to days. Such a behavior can be indicative of a larger number of clumps reaching the interaction cone when the pulsar was significantly far from periastron.
@@
Interestingly, the Bayesian analysis revealed the presence of short timescale variability well before periastron at and days, with fluxes comparable to the spin-down luminosity. While the Fermi/LAT sensitivity does not allow us to firmly conclude on the exact flux level of these flaring events we note that pre-periastron flares with spin-down-exceeding luminosity are hard to interpret within the discussed model. We also note that pre-periastron flares with a similar luminosity were not observed during the 2010-2021 periastron passages.
@@
The spectral behavior of the system in the GeV band differs from the one observed during the 2017-2021 periastron passages. Contrary to previous periastra, in 2024, PSR B125963 is detected with a higher significance before periastron ( to days, ) and a lower significance after periastron ( to days, ). The low significance of the detection during these time periods did not allow us to detect a spectral curvature contrary to previous passages. The flaring-period-averaged flux ( to days) is also somewhat lower than the one seen in 2017 (see Fig. 4 and Table 1).